### v2

# Training-Free Adaptation of New-Generation LLMs using Legacy Clinical Models

Abstract
Adapting language models to the clinical domain through continued pretraining and fine-tuning requires costly retraining for each new model generation. We propose Cross-Architecture Proxy Tuning (CAPT), a model-ensembling approach that enables training-free adaptation of state-of-the-art general-domain models using existing clinical models. CAPT supports models with disjoint vocabularies, leveraging contrastive decoding to selectively inject clinically relevant signals while preserving the general-domain model’s reasoning and fluency. On six clinical classification and text-generation tasks, CAPT with a new-generation general-domain model and an older-generation clinical model consistently outperforms both models individually and state-of-the-art ensembling approaches (average +17.6% over UniTE, +41.4% over proxy tuning across tasks). Through token-level analysis and physician case studies, we demonstrate that CAPT amplifies clinically actionable language, reduces context errors, and increases clinical specificity.
Training-Free Adaptation of New-Generation LLMs
using Legacy Clinical Models
Sasha Ronaghi1, Chloe Stanwyck1, Asad Aali2, Amir Ronaghi3, Miguel Fuentes4, Tina Hernandez-Boussard1,4, Emily Alsentzer1, 1Department of Biomedical Data Science, Stanford School of Medicine, Stanford University 2Department of Radiology, Stanford School of Medicine, Stanford University 3Department of Radiology, MemorialCare 4Department of Medicine, Division of Computational Medicine, Stanford School of Medicine, Stanford University Correspondence: sronaghi@stanford.edu
1 Introduction
Despite advances in general-domain language models, their application to clinical settings remains limited by hallucinations, omission of critical details, and failures in clinical reasoning (Lehman et al., 2023; Hager et al., 2024; Asgari et al., 2025). These shortcomings arise because pretraining corpora contain limited representations of clinical data, such as electronic health records, due to privacy constraints (Singhal et al., 2023a).
Domain adaptation techniques such as continued pretraining and finetuning address these challenges but require resource-intensive retraining for each new model generation (e.g., MedPalm 1 $\rightarrow$ MedPalm 2) (Singhal et al., 2023b). This creates a lag between advances in base model capabilities and their clinical applicability, with substantial computational resources reinvested in domain adaptation for each new architecture.
We explore model ensembling as a training-free approach to combine the advanced reasoning of state-of-the-art general-domain models with the domain knowledge of legacy clinical models. Prior works show that integrating decoding-time probability distributions of general-domain models with heterogeneous architectures can yield efficient performance gains Yao et al. (2025). However, these approaches assume largely overlapping model capabilities and have not been explored for combining models with distinct strengths, where information must be selectively integrated rather than uniformly aggregated. For example, we seek to incorporate the learned clinical knowledge of an older-generation clinical model without inheriting its limitations in instruction following, reasoning, or robustness due to its architecture or catastrophic forgetting (Kirkpatrick et al., 2017).
Contrastive decoding offers a mechanism for isolating a model’s strengths by choosing tokens that maximize the likelihood difference between expert and amateur models, amplifying the expert’s behavior (Li et al., 2023a). Proxy tuning leverages contrastive decoding to efficiently tune a large pretrained model by combining its logit distribution with the distributional delta between a smaller fine-tuned model and its untuned, base counterpart (Liu et al., 2024). However, proxy tuning assumes shared tokenization across models, restricting architectural diversity and preventing reuse of domain-adapted models built on older architectures.
Here, we propose Cross-Architecture Proxy Tuning (CAPT), a probability-level ensembling method that supports models with disjoint vocabularies and leverages contrastive decoding to selectively incorporate the specialized knowledge of a domain-adapted model. Our contributions include:
  • We introduce CAPT which enables reuse of legacy domain-adapted clinical models for training-free adaptation of newer-generation models. Across six clinical classification and text-generation benchmarks, CAPT consistently outperforms prior ensembling approaches, with an average improvement of 17.6% over UniTE and 41.4% over proxy tuning across metrics and tasks.
  • Through a token-level analysis, we illustrate that CAPT selectively integrates the clinical model’s knowledge for clinically actionable tokens, while the general model controls linguistic structure and formatting tokens.
  • In a physician-led case study, we demonstrate that CAPT-generated outputs contain more precise clinical terminology and context-appropriate recommendations.
1.1 Related Works
Model Ensembling. Probability-level model ensembling methods for heterogeneous architectures have primarily been explored with multiple general-domain models (Xu et al., 2024; Huang et al., 2024). UNiTE, the strongest-performing method to date, unions each model’s top-$k$ tokens at every decoding step and combines probabilities via re-tokenization (Yao et al., 2025). Prior analyses show ensembling methods only outperform the ensemble’s strongest model when performance gaps are within 10% (Yao et al., 2025). This limits applicability when combining new- and old-generation architectures where gaps can arise from architectural advances alone (Bedi et al., 2025). In contrast, CAPT is designed for models with asymmetric capabilities and enables selective integration of learned information from older domain models.
Contrastive Decoding. Contrastive decoding has been used to steer pretrained models toward desirable behaviors, including reduced toxicity (Liu et al., 2021), increased helpfulness (Mitchell et al., 2023), selective unlearning (Suriyakumar et al., 2025), and improved coding, QA, and math performance (Liu et al., 2024). These methods require shared vocabularies, and are evaluated in task- or instruction-tuning settings that are in-distribution of the large model’s training corpora. Our approach supports models with disjoint vocabularies and is evaluated on clinical domain adaptation, a substantially out-of-distribution setting (Kim et al., 2025).
2 Methods
Cross-Architecture Proxy Tuning (CAPT). We hypothesize that general-domain models encode substantial medical knowledge but lack exposure to clinical practice patterns and stylistic conventions reflected in clinical notes. Accordingly, the new-generation general-domain model should lead generation to preserve fluency and reasoning, while the clinical model selectively interjects for tokens associated with domain-specific reasoning or stylistic patterns. To achieve this, CAPT re-ranks the top-$k$ candidate tokens proposed by the general-domain model using a log-probability offset defined by the difference between the clinically trained model and its untrained base counterpart.
Formal definition. Let $M_{\text{new}}$ be a new-generation general-domain language model with tokenizer $\mathcal{T}_{\text{new}}$ and vocabulary $\mathcal{V}_{\text{new}}$. Let $M_{\text{old-clin}}$ denote an older-generation clinically trained model and $M_{\text{old}}$ its untrained base counterpart; both share tokenizer $\mathcal{T}_{\text{old}}$ and vocabulary $\mathcal{V}_{\text{old}}$.
Given a context $\mathbf{x}_{1:t}$, the models define next-token log-probabilities $\log p_{\text{new}}(\cdot\mid\mathbf{x}_{1:t})$ over $\mathcal{V}_{\text{new}}$ and $\log p_{\text{old-clin}}(\cdot\mid\mathbf{x}_{1:t})$, $\log p_{\text{old}}(\cdot\mid\mathbf{x}_{1:t})$ over $\mathcal{V}_{\text{old}}$. We restrict computation to the candidate set $\mathcal{C}_{t}=\text{Top-$k$}\big(\mkern-1.0mu\log p_{\text{new}}(\cdot\mid\mathbf{x}_{1:t})\big)\subseteq\mathcal{V}_{\text{new}}$, the $k$ most likely next-tokens under $M_{\text{new}}$.
Because the models operate over different vocabularies, we define a decode–retokenize mapping $f:\mathcal{V}_{\text{new}}\rightarrow\mathcal{V}_{\text{old}}$ that projects each candidate token from the new model into the clinical model’s vocabulary by decoding a token from $\mathcal{V}_{\text{new}}$ to its string form and re-tokenizing the string using $\mathcal{T}_{\text{old}}$. For $i\in\mathcal{C}_{t}$, we define $f(i)$ as the first non-space token in $\mathcal{T}_{\text{old}}(\mathrm{decode}_{\text{new}}(i))$. For each candidate token $i\in\mathcal{C}_{t}$, we update its log-probability:
$s(i)=\log p_{\text{new}}(i\mid\mathbf{x}_{1:t})\;+\\ \quad\alpha\Big(\!\log p_{\text{old-clin}}(f(i)\mid\mathbf{x}_{1:t})-\log p_{\text{old}}(f(i)\mid\mathbf{x}_{1:t})\!\Big)$
which adds a correction representing the isolated clinical-domain signal. The selected token, $i^{*}=\arg\max_{i\in\mathcal{C}_{t}}s(i)$, is appended to the context. We set $k=20$, reflecting the top-20 log-probabilities exposed by black-box API providers (OpenAI, 2025; Google DeepMind, 2025) and fix $\alpha=1.0$. See Appendix B for hyperparameter analysis.
Models and Baselines. We use Qwen3-30B (Yang et al., 2025) as the new-generation general-domain model ($M_{\text{new}}$) and MeLLaMA-13B-chat (Xie et al., 2024) as the old-generation clinical model ($M_{\text{old-clin}}$). MeLLaMA-13B-chat is continually pre-trained from LLaMA-2-13B-base ($M_{\text{old}}$) and instruction-tuned on clinical notes and tasks. We compare against proxy tuning (Liu et al., 2024) and UNiTE (Yao et al., 2025), the strongest probability-level ensembling method to date. Since proxy tuning requires a shared tokenizer between models, we use LLaMA-2-70B-chat ($M_{\text{old-L}}$) (Touvron et al., 2023) as its general-domain model.
Table 1: Performance comparison across tasks and methods. Bold indicates best performance. F1 = Macro-F1, Acc = Accuracy, LLM-J = MedHelm LLM-jury score (average out of 5), %RF = MedVAL % of risk-free outputs.
Classification Text-Generation
MedNLI MTS-Specialty CLIP MTS-Procedure MIMIC-RRS MIMIC-BHC
Method / Models F1 Acc F1 Acc F1 Acc LLM-J %RF LLM-J %RF LLM-J %RF
Old-gen large general ($M_{\text{old-L}}$) LLaMA-2-70B-chat 0.598 0.635 0.094 0.280 0.097 0.205 3.214 26.67 2.453 57.50 3.323 13.00
Old-gen clinical ($M_{\text{old-clin}}$) MeLLaMA-13B-chat 0.539 0.535 0.080 0.210 0.192 0.785 1.989 45.83 3.379 35.00 2.906 14.14
New-gen general ($M_{\text{new}}$) Qwen3-30B 0.858 0.850 0.140 0.350 0.240 0.540 3.825 63.33 4.394 65.00 3.932 30.00
Proxy Tuning $M_{\text{old-L}}$ + $M_{\text{old-clin}}$ + $M_{\text{old}}$ 0.575 0.620 0.095 0.275 0.143 0.420 3.466 55.00 3.898 34.17 3.456 8.00
UniTE $M_{\text{new}}$ + $M_{\text{old-clin}}$ 0.873 0.875 0.060 0.100 0.217 0.560 3.486 50.00 3.903 46.67 3.747 20.00
CAPT $M_{\text{new}}$ + $M_{\text{old-clin}}$ + $M_{\text{old}}$ 0.886 0.885 0.135 0.320 0.224 0.580 3.882 70.83 4.414 56.67 3.973 31.00
Evaluation Tasks. For classification, we report Macro-F1 and accuracy on 200 random test samples from: MedNLI (Romanov and Shivade, 2018), classifying premise-hypothesis pairs; MTS-Specialty (MTSamples, 2025), identifying medical specialties from transcriptions; and CLIP (Mullenbach et al., 2021), extracting physician action items from discharge notes. For text generation, we report MedHELM LLM-jury scores (Bedi et al., 2025) and MedVAL percent of risk-free outputs (Aali et al., 2025), both validated against clinician judgments, on 100 samples from: MIMIC-BHC (Aali et al., 2024), generating brief hospital course summaries; MTS-Procedures (MTSamples, 2025), generating treatment plans; and MIMIC-RRS (Chen et al., 2023), generating radiology study summaries. See Appendix C for details on metrics and tasks.
Token-level Analysis and Case Study. We categorized 280 tokens from the top 25% most frequently generated tokens in the MTS-Procedure task into semantic categories (Appendix D). For each category, we report the mean per-token log-probability difference between the clinical model $M_{\text{old-clin}}$ and its base counterpart $M_{\text{old}}$. To understand how CAPT’s token-level shifts affect clinical utility, two board-certified physicians annotated five MTS-Procedure outputs (Appendix E).
3 Results
CAPT outperforms baselines and existing model-ensembling approaches. Table 1 summarizes performance across classification and text generation tasks. $M_{\text{new}}$ outperforms $M_{\text{old-clin}}$ by an average of 53.6% in Macro-F1 on classification benchmarks and 52.6% in LLM-jury scores on text-generation benchmarks. This performance gap likely explains why UNiTE, which averages the probability scores of $M_{\text{old-clin}}$ and $M_{\text{new}}$, performs worse than $M_{\text{new}}$ alone by 8–28% on average. While proxy tuning narrows the gap between $M_{\text{old-L}}$ and $M_{\text{old-clin}}$, it is consistently outperformed by $M_{\text{new}}$, underscoring how architectural advances alone can substantially improve performance. CAPT outperforms $M_{\text{new}}$ on 8 of 12 metrics, while UniTE improves on only 2. Averaged across all tasks and metrics, CAPT outperforms UniTE and proxy tuning by 17.6% and 41.4%, respectively. For text-generation tasks, CAPT yields substantially more risk-free outputs (+35.85% vs. UniTE; +63.12% vs. proxy tuning), indicating improved clinical safety alongside performance gains.
CAPT selectively integrates clinical domain knowledge. Figure 1 summarizes mean log-probability shifts by semantic token category. Positive values indicate categories for which CAPT increases token preference via the clinical-model offset between $M_{\text{old-clin}}$ and $M_{\text{old}}$, while negative values indicate categories for which the offset decreases preference and $M_{\text{new}}$ retains stronger influence. $M_{\text{old-clin}}$ strongly influences tokens associated with clinical decision-making (e.g., Clinical Decision Action Headers, Clinical Assessment Terms) and documentation style (e.g., Clinical Hedging, Condition State Descriptors). In contrast, medical knowledge categories (e.g., Diagnostics, Medical Roots) exhibit relatively small shifts, consistent with CAPT primarily affecting how the knowledge of $M_{\text{new}}$ is expressed to match clinical documentation conventions rather than overriding its content. Linguistic structure and formatting remain dominated by $M_{\text{new}}$ (e.g., Formatting, General Morphemes, Medical Suffixes).
Figure 1: Mean log-probability offset between $M_{\text{old-clin}}$ and $M_{\text{old}}$ of generated tokens by semantic category. Positive shifts (orange) indicate increased influence of $M_{\text{old-clin}}$, while negative shifts (blue) of $M_{\text{new}}$. All token categories are shown in Appendix D, Figure 3.
Figure 2: CAPT-generated post-operative management plan for an MTS-Procedure example involving a forearm arteriovenous graft (connecting an artery to a vein). Orange and blue highlights indicate stronger influence from $M_{\text{old-clin}}$ and $M_{\text{new}}$, respectively. Bolded tokens denote when the top-choice token changed after CAPT adjustment, with the original top-choice crossed out. Green circles mark tokens discussed in main text. The full output is shown in Appendix E, Figure  4.
CAPT reduces context-specific errors and sharpens clinical terminology. A physician assessed outputs for accuracy, appropriateness, and overall utility. Figure 2 shows a representative sample of how CAPT’s token-level shifts improve clinical accuracy and specificity. For example, CAPT replaces “72” with “48” in the postoperative monitoring timeline, narrowing the range from 24-72 hours to 24-48 hours in line with typical discharge practices for this procedure. Similarly, $M_{\text{new}}$ recommends monitoring “cardiac” symptoms, which are not relevant to a forearm graft; CAPT correctly emphasizes perfusion (i.e. blood flow). CAPT also replaces $M_{\text{new}}$’s “neuro” token, which likely implies “neuro check,” with “dist.” This change directs generation towards monitoring “distal limb circulation,” which encompasses relevant components of the neurologic assessment (e.g., sensation and motor function) as well as other important checks (e.g., capillary refill). Finally, CAPT shifts generation toward more precise clinical terminology. For example, CAPT inserts “AV” before “graft” to clarify the graft type and replaces “agents” with “analgesics” to specify pain medication rather than any generic medication class.
A limitation is that CAPT can preserve over-explanatory content intended for non-expert audiences (e.g., enumerating infection signs or naming specific antibiotics), which are unnecessary for nurses responsible for post-operative management. These additions often appear in parentheses (e.g., “(”) that are favored by $M_{\text{new}}$, reflecting its general-audience training objective. Adaptively scaling the clinical-model adjustment coefficient (e.g., $\alpha$) could align outputs with user expertise. See Appendix E for extended analysis of Figure 2’s full output, LLM jury evaluation, comparison outputs of $M_{\text{new}}$, and additional examples.
4 Conclusion
We present CAPT, a model-ensembling strategy that enables clinical adaptation of newly-released models using legacy clinical models. We demonstrate how CAPT selectively integrates domain knowledge, improving performance, clinical utility, contextual accuracy, and lexical precision.
5 Limitations
Evaluation. Evaluating the clinical utility of open-ended LLM-generated outputs remains difficult without manual expert review, which is costly and does not scale. While classification tasks provide standardized, widely used benchmarks, clinical LLMs are more often deployed in text-generation tasks (Raji et al., 2025; Corporation, 2025; Shah Lab, 2025; Small et al., 2024), as smaller, simpler models can already achieve strong performance on basic classification tasks (Li et al., 2023b). In our text-generation tasks, we report peer-reviewed evaluation metrics: MedHELM LLM-jury scores assess accuracy, clarity, and completeness against gold-standard references (Bedi et al., 2025) and the percentage of risk-free outputs as determined by MedVAL, a fine-tuned clinical evaluation model (Aali et al., 2025). While both metrics have been benchmarked against physician reviews, they provide limited insight into the multi-dimensional nature of output quality. In practice, we observed that the MedHELM LLM jury reasoning for each score differed substantially from physician analysis, as the LLM jury did not highlight specific contextual accuracies or information.
To address this gap, we include token-level analyses and an expert qualitative case study. As demonstrated by the case study, tokens that most strongly affect clinical utility are often highly context-dependent and infrequent; consequently, while token-category analyses reveal broad trends, clinically meaningful improvements are difficult to capture through aggregate quantitative analyses alone. Future work should explore evaluation frameworks that better capture clinically meaningful improvements in open-ended generation.
Inference-time costs. Although CAPT does not introduce additional training costs, it increases inference-time computational overhead by requiring forward passes through three models. In our experiments, all three models were run on a single A100 GPU using 4-bit quantization. We did not observe a significant wall-clock latency difference between Qwen3-30b alone and CAPT (which additionally includes MeLLaMA-13b-chat and LLaMA-2-13-base), but CAPT requires greater computational resources. This may be a particular challenge for on-premise deployments at smaller health systems. In this study, we use MeLLaMA-13B models, which are the smallest models in the MeLLaMA family and are state-of-the-art among open-source clinical language models. Exploring CAPT with smaller or more specialized clinical models may further improve the feasibility of this approach.
6 Ethical Considerations
Our tasks are derived from two clinical datasets: MIMIC (Johnson et al., 2016) and MTSamples (MTSamples, 2025). MIMIC is a freely available, de-identified dataset of patients admitted to critical care units at Beth Israel Deaconess Medical Center. MTSamples is a publicly available, de-identified collection of medical transcription sample reports contributed by users for reference purposes. In terms of data consent, MIMIC data was collected under its respective IRB approval where the requirement for individual patient consent was waived. We access MIMIC under a data use agreement. MTSamples consists of voluntarily contributed data. Neither dataset requires Ethics Review Board approval for secondary analysis, as we perform in this study. We do not collect any additional data for this study which would require Ethics Review Board approval.
Although both datasets are de-identified, the risk of re-identification remains, meaning that private health information of an individual could be leaked. This study does not uniquely increase this risk as we propose a methodological innovation operating on existing data rather than introducing new data or aiming to uncover sensitive aspects of data.
Appendix A Additional Related Work
Clinical Domain Adaptation. Clinical models such as Gatortron, clinicalBERT, NYUTron, Me-LLaMA, and Clinical-T5 have demonstrated the effectiveness of continued pretraining on large volumes of unlabeled medical text (Yang et al., 2022; Alsentzer et al., 2019; Jiang et al., 2023; Xie et al., 2024; Lehman and Johnson, 2023). Following continued pretraining, supervised fine-tuning can further improve performance on specific clinical tasks, instruction following, and alignment with human preferences (Han et al., 2025; Zhang et al., 2025; Singhal et al., 2023a; Wang et al., 2025; Zhang et al., 2023). While effective, these approaches are highly resource-intensive, especially continued pre-training: for example, Me-LLaMA conducted continued pretraining on LLaMA-2 base models using 129B tokens and 160 × 80 GB A100 GPUs. Our work complements these techniques by addressing their limited transferability, as the benefits of such costly adaptations are typically confined to the specific model on which they are performed.
Appendix B Hyperparameter Analysis
To evaluate the effect of $k$ and $\alpha$, we conducted experiments on two held-out datasets. We report Macro-F1 on the validation split of the MedNLI classification task, and MedHelm LLM-jury score on ACI-Bench, a text-generation task of generating a structured clinical note from patient-doctor dialogues (Yim et al., 2023). We use ACI-Bench because MedHelm does not provide a validation set for its text-generation tasks and we did not want to reduce sample size of our tasks. Table 2 shows that varying $k$ has minimal impact on performance. We select $k=20$ because black-box API providers provide the top-20 logprobs for each token generation (OpenAI, 2025; Google DeepMind, 2025). We set $\alpha=1.0$ because it achieves best performance on both tasks, as shown by Table 3.
$k$ MedNLI (Macro-F1) ACI-Bench (LLM-jury score)
5 0.8944 4.377
10 0.8944 4.380
15 0.8944 4.377
20 0.8992 4.379
Table 2: Top-$k$ parameter experiments.
$\alpha$ MedNLI (Macro-F1) ACI-Bench (LLM-jury score)
0.5 0.8486 4.360
0.7 0.8537 4.379
1.0 0.8619 4.621
Table 3: $\alpha$ parameter analysis.
Appendix C Experimental Setup
For classification tasks, we use prefix-constrained generation to constrain outputs to an explicit JSON schema consisting of a "reason" field (a free-text string with a maximum length of 600 characters) and a "label" field (a string or array of strings drawn from a predefined, task-specific label set). At each decoding step, the current prefix is used to determine the set of tokens that can legally follow while still permitting completion of a valid JSON object; all other tokens are masked out. In CAPT, the constraint is applied to the base model’s token distribution, since the base model alone determines the generated tokens. In proxy tuning and UNiTE, the constraint is applied to the token candidates considered during model combination, ensuring that only schema-valid tokens remain eligible after score aggregation. We use constrained decoding to reflect realistic deployment settings and to remove performance differences driven by instruction-following rather than underlying knowledge.
For text-generation tasks, we implement a custom model client within the MedHELM evaluation framework (Bedi et al., 2025). We use MedHELM because it provides a standardized, reproducible, and peer-reviewed evaluation pipeline across clinical tasks, and includes an LLM-jury metric for text-generation tasks which assesses on medical accuracy, completeness and clarity compared to the reference output. Through use of MedHELM, our results can be directly compared to public leaderboard results, and CAPT can be readily applied to any of the 37 MedHELM clinical classification, text generation, or question-answering tasks. In addition, we report the percentage of responses that received a Level 1 (risk-free) grade using MedVAL, a physician-aligned risk grading LLM-as-a-judge system (levels 1–4, where level 1 is risk-free and level 4 represents the highest level of clinical risk). This metric assesses the clinical utility of text-generation outputs at scale, in addition to our qualitative analyses.
Table 4 contains a description of each task and the prompt used. All models were 4-bit quantized to fit on NVIDIA A100 40GB GPUs.
Appendix D Token-Level Analysis
We manually categorized 280 tokens of the 1,455 the top 25% most frequently generated tokens in the MTS-Procedure task into semantic categories. The remaining tokens were excluded because they were ambiguous, incomplete subword fragments, or otherwise not semantically interpretable. All tokens had a minimum frequency of 6. Tables 5 and 6 show token categories and corresponding tokens. Because shifts are computed over generated tokens, positive values indicate tokens preferentially up-weighted by the clinical model, while negative values indicate tokens supported by the new-generation general-domain model, such that they remain selected even when the contrastive signal disfavors them. Figure 3 shows the full version of Figure 1, with all token categories shown. Due space constraints, Figure 1 only includes the categories discussed in the main text.
Appendix E Physician Case Study
We present five randomly selected full-length CAPT outputs from the MTS-Procedure task, annotated with token-level shifts and expert qualitative analysis. The analysis was conducted by two authors of this paper, an interventional radiologist with 25 years of experience writing post-operative management plans and an obstetric gynecologist with expertise in clinical language models. Both were provided verbal instructions. We use MTS-Procedure because it is publicly available; our other text-generation tasks rely on MIMIC (Johnson et al., 2016), whose raw data cannot be shared. Figures 471013, and 16 contain the annotated CAPT outputs with expert qualitative analysis. For each case, we provide the LLM jury evaluation (Figures 58111417), and the corresponding output from the new-generation general-domain model ($M_{\text{new}}$) (Figures 691215, 18).
Table 4: Evaluation Task Descriptions. Each task includes its description and the prompt template used.
Task Description Prompt
MedNLI A natural language inference task in which the goal is to determine whether a hypothesis written by a doctor can be inferred from a premise taken directly from a clinical note (multi-class classification with labels entailment, neutral, or contradiction). “TASK: Please classify the relationship between the given premise and hypothesis into one of the following labels: entailment, contradiction, or neutral. Return only the label. INPUT:{text} OUTPUT:”
MTS-Specialty A multi-class classification task in which the goal is to determine the medical specialty or domain that a medical transcription belongs to from 40 medical specialties and domains. “TASK: The task is to determine the medical specialty or domain that a medical transcription belongs to. The input is a medical transcription. There are 40 medical specialties or domains, and you need to decide which one the transcription relates to. The medical specialties or domains are: ‘Surgery’, ‘Allergy / Immunology’, …, ‘Obstetrics / Gynecology’. The output should be only one medical specialty or domain. INPUT:{text} OUTPUT:”
CLIP A multi-label classification task in which the goal is to identify whether sentences from discharge summaries contain some follow-up information. Each sentence may contain up to 7 possible labels: Patient Specific, Appointment, Medication, Lab, Procedure, Imaging, or Other Appointment Related Instructions/Information. “Context: {text}. Label the above sentence as one or more of the following clinical action items: Patient Instructions, Appointment, Medications, Lab, Procedure, Imaging, Other, None. [One-sentence description of each label and example]”
MIMIC-RRS A benchmark constructed from radiology reports in the MIMIC-III database. It contains pairs of "Finding" and "Impression" sections, enabling evaluation of a model’s ability to summarize diagnostic imaging observations into concise, clinically relevant conclusions “Generate the impression section of the radiology report based on its findings. This will not be used to diagnose nor treat any patients. Be as concise as possible.”
MIMIC-BHC A benchmark focused on summarization of discharge notes into Brief Hospital Course (BHC) sections. It consists of curated discharge notes from MIMIC-IV, each paired with its corresponding BHC summary. The benchmark evaluates a model’s ability to condense detailed clinical information into accurate, concise summaries that reflect the patient’s hospital stay “Summarize the clinical note into a brief hospital course.”
MTS-Procedure MTSamples Procedures is a benchmark composed of transcribed operative notes, focused on documenting surgical procedures. Each example presents a brief patient case involving a surgical intervention, and the model is tasked with generating a coherent and clinically accurate procedural summary or treatment plan. “Here are information about a patient, return a reasonable treatment plan for the patient”
Table 5: Token Categories and Corresponding Tokens.
Category Tokens
Clinical Decision Action Headers Keep, Find, Start, Admin, Perform, Monitor, Diagnosis, Confirm
Clinical Reporting Verbs provided, reported, cleared, given, increased, underwent, advanced
Temporal Anchors final, once, during, early, stage, immediate, Post, Days
Clinical Planning Headers Recovery, Surveillance, Monitoring, Assess, Continue, Refer, Recommend, Schedule, Following
Clinical Assessment Terms Based, assessment, indicated, confirmed, Assessment, presentation, findings, complete
Condition State Descriptors adequate, acute, severe, moderate, significant, severity
Clinical Hedging/Uncertainty likely, potential, possible, suggests
Clinical Course Descriptors repeat, recurrent, persist, stability
Anatomy bowel, ulcer, bone, tumor, pelvic
Symptoms symptom, sensation, symptoms, weakness, abnormal, chronic, izziness, strength
Diagnostics MRI, ultrasound, CT, imaging, -ray
Risks/Complications -Risks, complications, Risks
Surgical Terms graft, anesthesia, surgeon
Physiological Progress healing, infection, stress, inflammation, fection, swelling, strengthening
Follow-up Care regimen, instructions, Avoid, future, post, exercises, referral, specialist, appointment
Diagnoses pneumonia, carcinoma, fever, hemorrh
Medical Roots bron, fib, metast, throm, neph, Hem, uro
Time Units activities, times, count, Week, months, minutes, hours, daily, days
Treatment Terms rehabilitation, antibiotics, steroid, course
Time Demographic Modifiers -Year, -term, -old, -year
Reproductive and Gynecologic Terms vaginal, cervical, ovarian, uterus, pregnancy, fetal, Breast, breast
Gender Terms Female, female, Male, male, men
Musculoskeletal Function joint, injury, mobil, tear, Motion, motion, activity
Formatting Tokens (, :***, ####, :*, ###, ").", ), ":", **, ., **, —, *
Table 6: Token Categories and Tokens (Continued).
Category Tokens
General Morphemes ism, able, ised, riage, ization, oph, omet, aging, col, vic, ions, olin, ping, ural, bar, um, ute, cin, ized, aneous, ility, orb, verse, ore, icture, ister, situ, atin, ops, cess, ency, par, fen, ond, ges, ab, ence, urrent, ating, ch, ot, ated, es, sub, en, ach, us, ef, ist, air, ox, com, uous, ps, ob, ing, th, uff, el, pol, et, ain, uture, tes, atory, ification, aph, gest, igation, ol, actic, osis, ose, ologist, ology, ogram
Medical Suffixes oglobin, urgery, emia, monary, inine, ascular, iated, axis, ement, adder, ication, isc, odal, opsy, otherapy, astro, ycin, scopic, umor, ngthen, ipation, struct, onic, opathy, urgical, ateral, omat, oid, ortic, ricular, rosis, esthesia, opic, oster, operative, olic, ard, ulent, icated, yps, agnosis, ologic, iotic, otomy, opath, oscopic, ausal, dic
Figure 3: Full version of Figure 1 with all token categories shown. Mean log-probability offset between $M_{\text{old-clin}}$ and $M_{\text{old}}$ of generated tokens by semantic category. Positive shifts (orange) indicate increased influence of $M_{\text{old-clin}}$, while negative shifts (blue) of $M_{\text{new}}$.
Figure 4: Case Study 1 - CAPT Output, LLM Jury Score = 4.22. Majority of comments related to this output can be found in Section 3, Results. We include additional comments here. CAPT increases clinical specificity, resulting in a more clinically actionable and descriptive plan. CAPT heavily preferences “proph” after “Administer,” emphasizing the preventative goal of the recommended antibiotic administration. CAPT also replaces “institutional” with “surgical” before “protocol,” increasing specificity as typically each institution has a protocol for each surgery. CAPT replaces the generic “clinical exam,” which can include physical and imaging exams, with “physical exam”, which is more specific to the type of follow-up care the patient needs. CAPT also adds monitoring “thrill,” which is an important physical exam function indicating that the graft is functioning. CAPT also uses more clinically relevant terminology. For example, CAPT replaces “optimize” with “resume” when discussing hemodialysis, which is an important distinction as it is not possible to optimize and that language is not used clinically. Additionally, CAPT describes a “native” graft, which is the most common and best way of performing the graft. Despite these improvements, there remain limitations. For example, the plan advises "If the patient is not yet on dialysis, initiate or resume hemodialysis" and to "Assess current dialysis adequacy," which does not provide any patient-specific advice and would not be useful language in a treatment plan. Similarly, the advice for "strict glycemic control (if diabetic)" and to "Continue antihypertensive agents preferentially those safe in CKD" are also not patient-specific and are far more general than would be found in a real patient treatment plan. This is likely an artifact of the task itself, where insufficient patient information (e.g. about prior medications or medical history) is given to provide detailed patient-specific advice, though is worth noting as a weakness.
Figure 5: Case Study 1 - CAPT Output, LLM Jury Evaluation, LLM Jury Score = 4.22.
Figure 6: Case Study 1 - $M_{\text{new}}$ Output, LLM Jury Score = 4.11.
Figure 7: Case Study 2, LLM Jury Score = 4.44. CAPT adds clinically specific details which make the output more realistic and accurate. For example, CAPT adds “pain control,” a top priority as the epidural analgesia wears off, and heavily preferences “neurological status,” which should be specifically monitored after epidural. CAPT also replaces “bladder” with “urinary” output, which more clinically realistic phrasing. CAPT adds “breakthrough” in the sentence about short-acting opioids for breakthrough pain, which is clinically more realistic – post-c-section pain control would rely on scheduled Tylenol (and usually NSAIDs) plus as-needed opioid for breakthrough pain. In wound care, $M_{\text{new}}$ prefers swelling, which is more of a general sign of infection, which in c-sections particularly, will be less apparent because the surgery is abdominal. CAPT preferences towards erythema, discharge, and dehiscence, which are more procedure-specific. CAPT replaces “infection” with “thromboembolism,” which is more relevant to ambulation as early ambulation reduces risk of thromboembolism but does not reduce risk of infection. CAPT adds “fundal” height, a very procedure-specific sign to monitor. CAPT replaces “foul” with “sudden,” reflecting a shift in the focus from monitoring for signs of infection (could be accompanied by foul odor of lochia, but that’s not a primary sign) to postpartum hemorrhage (sudden increase in lochia could signify a delayed postpartum hemorrhage). Similarly, CAPT replaces “fever” with “passage of large clots,” adding specific warning signs for delayed postpartum hemorrhage. In the follow-up paragraph, CAPT adds the word “contraception” which would be a key part of the follow-up visit (usually 6 week) and replaced “safety” with “breastfeeding compatibility,” which is more accurate. One limitation is that the instructions say to avoid NSAIDs, which is not unreasonable, but would be atypical to avoid unless there had been a huge hemorrhage and 1000 mL is relatively high, but not unusually high. Additionally, $M_{\text{new}}$ preferences over-explanatory generation (e.g., defining opioids as oxycodone, hydromorphone), which is not necessary for the clinical setting.
Figure 8: Case Study 2 - CAPT Output, MedHELM LLM Jury Evaluation, LLM Jury Score = 4.44.
Figure 9: Case Study 2 - $M_{\text{new}}$ Output, LLM Jury Score = 4.66.
Figure 10: Case Study 3, LLM Jury Score = 4.77. CAPT adds crucial contextual details. CAPT replaces “Diagnostic” with “Dietary,” which is the most relevant aspect of the plan besides immediate care, because the procedure is related to the esophageal tube. In describing when to administer the prokinetic agent, CAPT adds that it should be before meals, which is a crucial detail as the medication is meant to help with gastric mobility. CAPT also adds that the patient should have a follow up endoscopy, which is necessary if a patient worsens after initially improving. CAPT enhances detail. For example, it adds “prophylactically,” which clarifies that the suggestion of antibiotics would be preventative, in the absence of active infection. Despite these improvements, the output contains some inaccuracies. For example, it suggests performing a barium esophagram, which would not be appropriate if the physician is concerned about perforation, as they should instead perform a water soluble esophagram. Another limitation of this output is that it reads less like a specific patient treatment plan, and more like a protocol for the entire hospital. For example, the plan suggests keeping the patient NPO (nil per os, or nothing to eat) until esophageal integrity is confirmed, and performing additional testing “if there is any clinical concern for perforation.” In a specific patient plan, recommendations would be based on the levels of concern for perforation in this patient, and a definitive recommendation would likely be made. This vagueness likely stems from the prompt not including enough patient details.
Figure 11: Case Study 3 - CAPT Output, MedHELM LLM Jury Evaluation, LLM Jury Score = 4.77.
Figure 12: Case Study 3 - $M_{\text{new}}$ Output, LLM Jury Score = 4.77.
Figure 13: Case Study 4, LLM Jury Score = 4.44. CAPT heavily favors “Mechanical”, which specifies that the patient is intubated. CAPT replaces “patient” with “INR,” which is a specific measure from a blood test that should be monitored closely. CAPT adds “(mechanical valve indication),” which justifies the INR range in a way that would be helpful for a clinician reading the note and is realistic in terms of how notes are usually written. CAPT also replaces “rate” with “beta” (for beta-blockers), which is the specific class of medications which slow heart rate. CAPT also replaces “cornerstone,” a word not typically used in real-world treatment plans and more often when describing institutional protocols or in academic settings, with “treatment.” This highlights how the general-domain model’s training corpora of publicly available literature can result in outputs that do not reflect real-world clinical note style, and the ability of CAPT to address this tendency. The output still contains limitations: It is over-explanatory (e.g.) with descriptions of electrolytes that need to be monitored), highlighting shortcomings of CAPT in changing highly confident general-domain tokens (e.g., parantheses) which result in redundancies. Additionally, the output contains unrealistic information. For example, the output says that glucose targets should be between 110-150, and it is normally 140-180. The INR target for mechanical valves is usually 2.0-3.0, not 2.5-3.5. INR would likely be monitored more often - often daily as inpatient, then often 1-2x/week initially and 3-6 months later on even if stable. Epidural is not usually used post-sternotomy because anticoagulation is very important for these patients (including this one), so epidural placement is generally unsafe Beta-blockers aren’t used for afterload reduction, and more commonly used for rate control. These inconsistencies highlight that there is room for improvement with the clinical model to ensure safer CAPT outputs.
Figure 14: Case Study 4 - CAPT Output, MedHELM LLM Jury Evaluation, LLM Jury Score = 4.44.
Figure 15: Case Study 4 - $M_{\text{new}}$ Output, LLM Jury Score = 4.11.
Figure 16: Case Study 5, LLM Jury Score = 4. CAPT increases specificity and accuracy. CAPT adds “bilateral” and changes “lymph node” to “lymph nodes,” which is more precise phrasing given that the procedure involved bilateral inguinal lymphadenectomy. For close monitoring, CAPT also changes from “late radiation” to “late complications,” which is the accurate wording for this context. CAPT replaces “pall” with “plastic”; this was likely the beginning of “palliative,” which is less appropriate because plastic surgery is a likely next step after this procedure. CAPT changes “including inspection” to be more specific: “including vaginal and inguinal region assessment.” CAPT replaces “manage lymphedema prophylaxis” with “prophylactically with compression garments and education,” which is more correct as one wouldn’t manage lymphedema prophylaxis. CAPT replaces “Provide written information and resources” with more specific instructions of “Provide written information and involve the patient in treatment decisions.” In the summary, CAPT adds “clinical” before “stage III,” which is more precise and reflective of clinical note style. Overall, these instructions are vague, likely because the prompt doesn’t contain much patient details.
Figure 17: Case Study 5 - CAPT Output, MedHELM LLM Jury Evaluation, LLM Jury Score = 4.
Figure 18: Case Study 5 - $M_{\text{new}}$ Output, LLM Jury Score = 4.44.